Three people pull simultaneously on a stubborn donkey. Jack pulls directly ahead of the donkey with a force of 63.3 N, Jill pulls with 71.9 N in a direction 45 to the left, and Jane pulls in a direction 45° to the right with 127 N. (Since the donkey is involved with such uncoordinated people, who can blame it for being stubborn?) Find the magnitude of the net force the people exert on the donkey Number What is the direction of the net with a positive sign for angles to the left and a negative sign for angles to the right force? Express this as the angle from straight ahead between 0 and 90 Number
Answer #1

given

F1 = 63.3 i N

F2 = 71.9 x ( - cos45 j + sin45 i )

F3 = 127 x ( cos45 j + sin45 i )

Fnet = F1 + F2 + F3

= ( 63.3 + 71.9 x sin45 + 127 x sin45 ) i + ( - 71.9 x cos45 + 127 x cos45 ) j

= 203.943 i + 38.96 j

then the magnitude is |Fnet| = ( 203.9432 + 38.962 )2

|Fnet| = 207.63 N

and the direction is = tan-1 ( 38.96 / 203.943 )

= 10.815o